# Multi-qubit Dynamical Decoupling for Enhanced Crosstalk Suppression

### v2

@@ I Introduction
The inherent noise in near-term quantum computers hinders the demonstration of practical quantum advantage. Recently, a variety of quantum error mitigation techniques have been developed to tackle this challenge, such as zero noise extrapolation (ZNE) [1, 2], probabilistic error cancellation (PEC) [3], and clifford data regression (CDR) [4], which require substantial overhead in quantum resources (such as the number of experiments executed) in order to model the noise and mitigate the resulting error. Quantum error suppression techniques, such as randomized compiling [5], pulse-scaling for Trotterized circuits [6], and dynamical decoupling (DD) [7, 8, 9], offer a reduction in error with no additional quantum resources. DD aims to reduce the decoherence error by decoupling the interaction between the quantum system and the environment. This process involves inserting a sequence of gates that compose the identity during the idle period when the qubit is not being used for computation. This entails very little classical overhead, as it only requires identifying the idle times and inserting sequences with a pattern of delays between the gates. It makes DD one of the simplest error suppression methods. Numerous DD sequences have been introduced over decades [10, 8, 11, 12, 13], leading to performance improvements on various modalities of quantum computers, including superconducting circuits [14, 15, 16, 17, 18, 19], trapped ions [20, 21], and neutral atoms [22]. These studies typically focus on a single qubit and its environment but quantum computers consist of many qubits coupled to each other. Notable exceptions [23, 18] explore static $ZZ$ crosstalk in multi-qubit systems. They propose dynamical decoupling (DD) protocols to protect the quantum systems from crosstalk and enhance the gate fidelity, similar in pattern to those we consider.
@@ II Theoretical analysis
$H_{SB}^{\prime}$ $=$ $(\sigma^{x}\otimes I_{SB})H_{SB}(\sigma^{x}\otimes I_{SB})$ (4)
so that the free evolution during $f_{\tau}$ is
$U_{SB}^{\prime}=e^{-it(\sigma^{x}\otimes B_{x}-\sigma^{y}\otimes B_{y}-\sigma^{z}\otimes B_{z})}.$ (5)
@@
where explicit references to the free evolutions $f_{\tau}$ have been omitted for simplicity. Noisy simulations illustrating the cancellation of $ZZ$ crosstalk via the use of staggered DD are shown in Appendix C.
In the following, we study the differences between staggered and standard DD sequences of different orders (X2, XY4, XY8) by inserting idle times (representing free evolution of time $2\tau$) between Clifford gates in two-qubit randomized benchmarking (RB) sequences [29, 30]. These RB sequences are performed individually and in conjunction with operations on an adjacent pair of qubits, and show that staggered DD consistently suppresses coherent error better both when it is caused by static transmon-transmon higher-level coupling and also cross resonance-driven crosstalk.
@@ III.1 Experimental Protocol
We select a subset of nine qubits from the 27-qubit ibm_cairo backend, which includes controlled-NOT (CX) pairs generated by both echoed cross resonance (ECR) [31, 32] and direct CX (DCX) [26]. We focus on physically connected qubit pairs due to their increased susceptibility to crosstalk. In this configuration, each pair neighbors the other, where operations on one pair can induce crosstalk in the adjacent pair [33, 34, 35]. We present the RB experiments in two experimental settings: (i) idle-idle in which the two pairs of qubits are simultaneously idle during their simultaneous RB sequences (aside from the relevant DD sequences) in Section III.2, and (ii) idle-driven in which the fidelity of the vulnerable pair of qubits is measured after the RB sequence (with varying DD sequence) while the other pair of qubits is driven by a random sequence of two-qubit gates (SU(4)) in Section III.3. The two-qubit individual RB circuit sets a baseline for the two experiments for comparison, where the $ZZ$ interaction between qubits in the pair is suppressed by calibrating rotary echo pulses on the target qubits [36]. This is compared to the same RB circuit but while a neighboring qubit pair is active, effectively exposing the RB circuit to static and/or driven crosstalk (the effect of static crosstalk in which the neighboring qubits are in the ground state has been calibrated out by the backend). The idle-idle and idle-driven experiments expand to involve two qubit pairs, totaling four qubits. Standard and staggered DD sequences are inserted during the idle periods. Specifically, in the idle-idle experiment, DD are applied to both qubit pairs, whereas in the second experiment, DD is only applied to the first qubit pair with RB sequences. These experiments highlight the difference in performance between standard and staggered DD sequences when the qubits are exposed to both stochastic dephasing and coherent crosstalk arising from static $ZZ$ coupling and cross-resonance driven $ZZ$. The quantum circuits were generated and executed using the Qiskit Experiments framework [37].
@@
We choose a pair of qubits $(q_{11},q_{14})$ and $(q_{12},q_{13})$ for study, that are particularly vulnerable to crosstalk from operations on neighboring qubit pair, as shown by the fidelity differences between figure 3a and figure 3b. For idle-idle experiment, we apply exactly the same RB circuit simultaneously to the adjacent pair of qubits. While the duration of the idle period remains the same for both qubit pairs, the starting points of these idle periods differ slightly due to the differences in duration of the two neighboring CX pairs. We use the As Late As Possible policy to schedule the circuits. For the idle-driven experiment, while one qubit pair undergoes the RB circuit, its adjacent pair is subjected to a circuit composed of different random SU(4) blocks, which continuously drives crosstalk during the RB experiment, as shown in figure 2e. The number of SU(4)s in the circuit is determined to ensure it has the same duration as the RB circuit. This experimental arrangement ensures the simultaneous execution of both circuits. Since no idle time is inserted between SU(4)s, during the idle periods of the RB circuit, the neighboring qubits in the SU(4) circuit remain active thereby contributing to driven crosstalk.
@@
In the two experiments, the CX gate is aligned with the native direction of either ECR or DCX gates, determined in calibration to that of higher fidelity. For the two adjacent qubit pairs, the neighbor qubits from each pair can be control-control, control-target, or target-target in their respective pairs. The performances across these different configurations, in terms of the type of CX gate and the type of connected neighbor qubits, are detailed in Appendix A.
@@ III.2 Idle-idle Experiment - Simultaneous Randomized Benchmarking
The RB and simultaneous RB (SRB) results for the qubit pair ($q_{11},q_{14}$) are illustrated in figure 3, showcasing the comparison between various X2 sequences (X2pm, X2pm-stag, X2pm-stag-inv) against the free evolution. In the SRB experiment, it involves the adjacent qubit pair of ($q_{11},q_{14}$), which is ($q_{12}$, $q_{13}$), with $q_{14}$ and $q_{13}$ being the target qubits in their respective pairs and neighboring each other. The figure shows the connectivity of the two qubit pairs, detailing the type of CX gate used (ECR or DCX) and identifying whether a qubit acts as a control (C) or target (T) for the neighboring qubit connections between pairs. The connection between the qubit pair ($q_{11},q_{14}$) is highlighted in red to indicate the specific results corresponding to this pair in the figure. The insertion of default X2 pulses results in unexpected behavior, even degrading the performance below that of the free evolution. The reason for this outcome is elaborated in Appendix  B.
We also demonstrate the results for XY4 and XY8 sequences in figure 4. Given the negligible difference between the results for staggered DD and its inverse, only the staggered DD is depicted. We observe “dropouts” in fidelity for a short idle times in the XY8 sequence which can be explained by constructive interference of higher frequency spectral noise components [38, 39]. In addition, the XY8 sequence shows consistently worse results than XY4. Therefore, our subsequent discussions will focus primarily on the X2 and XY4 sequences.
Figure 3: Idle-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with different DD sequences: X2pm, staggered X2pm, and inverse of staggered X2pm. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when SRB is applied to both qubit pairs. The significant difference in fidelity during free evolution between the RB and SRB experiments indicates additional crosstalk generated by the neighboring qubit pair.
Figure 4: Idle-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with XY4pm, staggered XY4pm, XY8pm, and staggered XY8pm. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when SRB is applied to both qubit pairs.
@@ III.3 Driven-idle Experiment - Randomized Benchmarking and Random SU(4)
Figure 5: Driven-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when only RB is applied. (b) The same probability when of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair while SU(4) is applied to its adjacent qubit pair $(q_{12},q_{13})$.
@@ IV Conclusion
Our proposed staggered DD protocol can be considered as an entry point for integrating DD sequences into multi-qubit quantum systems. Throughout the experiments, we have tailored our staggered DD protocol to fit X2, XY4, and XY8 sequences, incorporating them during unified idle times. Several advanced DD sequences, such as UDD [12] and UR [13], offering opportunities for further exploration. One natural following step is to apply the staggered DD protocol to these sequences and delve into DD scheduling methods for non-unified idle time in large quantum systems. Recently, many efforts are focused on understanding noise models for quantum hardware and developing corresponding error suppression or mitigation techniques [3]. An interesting direction is to synergize the staggered DD protocol with these techniques to enhance circuit fidelity further. Moreover, standard DD sequences have been employed to protect idle qubits for quantum error correction or detection codes [41, 42, 15, 43]. Incorporating our staggered DD protocol into quantum error correction codes offers a pathway toward achieving fault-tolerant quantum computing. Despite of the positive impact of DD sequences, challenges such as dropouts or instability across different qubits persist. Additional experiments are necessary to deepen the understanding of hardware noise sources and to devise more hardware-oriented DD sequences.
@@ Acknowledgements.
The authors acknowledge the IBM Quantum Network Hub at QuantUM Initiative of the Region Occitanie and University of Montpellier for providing additional quantum resources through its partnership with IBM. The authors would like to thank Bibek Pokharel for a careful reading of the manuscript and Ali Javadi-Abhari and Petar Jurcevic for insightful discussions.
@@ Appendix A Various Experimental Configurations
There are two variants of CX gates in IBM quantum hardware: Echoed Cross-Resonance CX (ECR), and Direct CX (DCX). We select several qubit pairs from ibmq_cairo including both ECR and DCX types, and list their respective amplitudes, duration, and error rates in table 2. Note that, CX gate is natively uni-direction in IBM quantum hardware, and the opposite direction is realized by adding extra single qubit gates. Therefore, we show explicitly the control and target qubits for each qubit pair in the table. Generally, DCX gate exhibits a lower error rate and shorter duration compared with ECR gate. However, DCX gate is more susceptible to crosstalk, as illustrated in figure 6. Here, the CNOT error rate, derived from SRB, exhibits a notable increase compared to the error derived from independent RB for DCX, whereas the CNOT error rate for the ECR only experiences a slight increase.
Figure 6: We perform RB and SRB on two qubit pairs and calculate the corresponding CNOT error rate. The selected qubit pairs employ either ECR or DCX type of CNOT gate. An increase in the CNOT error rate observed during SRB as compared to RB indicates the impact of crosstalk.
@@
Table 2: CX information.
CX(control-target) Type Amplitude Duration (dt) Error rate (%)
(10, 12) DCX 0.42 1008 0.66
(13, 14) DCX 0.33 992 0.49
(11, 14) DCX 0.49 848 1.37
(1, 0) ECR 0.91 2400 2.57
(4, 7) ECR 0.58 1184 1.64
(15, 12) ECR 0.52 2208 0.9
(12, 13) ECR 0.92 1760 1.4
@@
are shown in (a) and (b), on $[(q_{10},q_{12})-(q_{13},q_{14})]$ are shown in (c) and (d), on $[(q_{12},q_{13})-(q_{11},q_{14})]$ are shown in (e) and (f), and on $[(q_{15},q_{12})-(q_{13},q_{14})]$ are shown in (g) and (h).
Appendix B X2
The exploration of circuit fidelity for both idle-idle and driven-idle experiments reveal that the insertion of $X2$ sequence leads to worse outcomes than the baseline, as shown in figure 8, which is potentially due to the imperfections in the X gate. However, this was not reproducible in simulations with an error model informed by quantum process tomography. Rather, we suspect an observation similar to [17] performed on IBM quantum hardware, where fidelity decreases due to the asymmetry of the trajectory around the Bloch sphere during qubit rotations. Since an X gate corresponds to a $\pi$ rotation around the $x$-axis in one direction, this asymmetry accumulates after two consecutive X gates, thereby resulting in poorer performance compared to the baseline where no DD sequence is inserted. On the other hand, the Xp and Xm operations, which correspond to a $\pi$ rotation around the $X$ axis in opposite directions, effectively mitigate this asymmetry.

### v3

@@ I Introduction
The inherent noise in near-term quantum computers hinders the demonstration of practical quantum advantage. Recently, a variety of quantum error mitigation techniques have been developed to tackle this challenge, such as zero noise extrapolation (ZNE) [1, 2], Clifford data regression (CDR) [3], and probabilistic error cancellation (PEC) [4], which require substantial overhead in quantum resources (such as sampling overhead) in order to model the noise and mitigate the resulting error. Specifically, ZNE, CDR and PEC incur linear, polynomial, and exponential sampling overheads, respectively. Among these methods, only PEC provides an accurate and bias-free estimate. In contrast, quantum error suppression techniques, such as randomized compiling [5], pulse-scaling for Trotterized circuits [6, 7], and dynamical decoupling (DD) [8, 9, 10], offer a reduction in error with no additional quantum resources. However, randomized compiling necessitates running circuits multiple times and averaging the results for effective error suppression, which may introduce latency due to the increased bandwidth when implementing on cloud-based quantum computers. Pulse-scaling does not require additional resources if the default pulses are parametrically scaled, otherwise the overhead of additional calibration experiments must be considered. DD aims to reduce the decoherence error by decoupling the interaction between the quantum system and the environment. This process involves inserting a sequence of gates that compose the identity during the idle period when the qubit is not being used for computation. This entails very little classical overhead, as it only requires identifying the idle times and inserting sequences with a pattern of delays between the gates. It makes DD one of the simplest error suppression methods, as it involves no additional sampling or experiment execution overhead. Numerous DD sequences have been introduced over decades [11, 9, 12, 13, 14], leading to performance improvements on various modalities of quantum computers, including superconducting circuits [15, 16, 17, 18, 19, 20], trapped ions [21, 22], neutral atoms [23], and semiconductors [24]. DD achieves these performance improvements by reducing not only decoherence error but also crosstalk, which is a major error source in these quantum technologies [25, 26, 27, 28]. These studies typically focus on a single qubit and its environment but quantum computers consist of many qubits coupled to each other. Notable exceptions [29, 19, 30, 31] explore static $ZZ$ crosstalk in multi-qubit systems. They propose dynamical decoupling (DD) protocols to protect the quantum systems from crosstalk and enhance the gate fidelity, similar in pattern to those we consider.
@@ II Theoretical analysis
$H_{SB}^{\prime}$ $=$ $(\sigma^{x}\otimes I_{B})H_{SB}(\sigma^{x}\otimes I_{B}),$ (4)
where $I_{B}$ represents the identity operator on the bath, so that the free evolution during $f_{\tau}$ is
$U_{SB}^{\prime}=e^{-i\tau(\sigma^{x}\otimes B_{x}-\sigma^{y}\otimes B_{y}-\sigma^{z}\otimes B_{z})}.$ (5)
@@
where explicit references to the free evolutions $f_{\tau}$ have been omitted for simplicity. Noisy simulations illustrating the cancellation of $ZZ$ crosstalk via the use of staggered DD are shown in C. In the following, we study the differences between staggered and standard DD sequences of different orders (X2, XY4, XY8) by inserting idle times (representing free evolution of time $2\tau$) between Clifford gates in two-qubit randomized benchmarking (RB) sequences [37, 38]. These RB sequences are performed individually and in conjunction with operations on an adjacent pair of qubits, demonstrating that staggered DD consistently suppresses coherent error more effectively, whether it is caused by static transmon-transmon higher-level coupling or cross-resonance-driven crosstalk.
Figure 2: Qubit coupling graph of 27-qubit ibm_cairo backend. The selected qubits for experiments are highlighted in blue circles. The qubit connections follow the native directions. ECR and DCX types of CX gates are represented by yellow and purple colors respectively.
@@ III.1 Experimental Protocol
We select a subset of ten qubits from the 27-qubit ibm_cairo backend (the coupling graph is shown in figure 2), which includes controlled-NOT CX pairs generated by both echoed cross resonance (ECR) [39, 26] and direct CX (DCX) [34]. We conduct experiments on two qubit pairs to study the crosstalk between them. The two qubit pairs are chosen to be physically connected for two reasons: (1) Connected qubits within each pair can introduce static crosstalk. (2) Operations on one qubit pair can induce more driven crosstalk to its neighboring qubit pair [40, 41, 42]. We present the RB experiments in two experimental settings: (i) idle-idle in which the two pairs of qubits are simultaneously idle during their simultaneous RB sequences (aside from the relevant DD sequences) in Section III.2, and (ii) idle-driven in which the fidelity of the vulnerable pair of qubits is measured after the RB sequence (with varying DD sequence) while the other pair of qubits is driven by a random sequence of two-qubit gates (SU(4)) in Section III.3. The two-qubit isolated RB circuit sets a baseline for the two experiments for comparison, where the $ZZ$ interaction between qubits in the pair is suppressed by calibrating rotary echo pulses on the target qubits [43]. This is compared to the same RB circuit but while a neighboring qubit pair is active, effectively exposing the RB circuit to static and/or driven crosstalk (the effect of static crosstalk in which the neighboring qubits are in the ground state has been calibrated out by the backend). The idle-idle and idle-driven experiments expand to involve two qubit pairs, totaling four qubits. Standard and staggered DD sequences are inserted during the idle periods. Specifically, in the idle-idle experiment, DD are applied to both qubit pairs, whereas in the second experiment, DD is only applied to the first qubit pair with RB sequences. These experiments highlight the difference in performance between standard and staggered DD sequences when the qubits are exposed to both stochastic dephasing and coherent crosstalk arising from static $ZZ$ coupling and cross-resonance driven $ZZ$. The quantum circuits were generated and executed using the Qiskit Experiments framework [44].
@@
We choose two qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ for study, that are particularly vulnerable to crosstalk from operations on neighboring qubit pair and demonstrates stable output, as shown by the fidelity differences between figure 4a and figure 4b. Data for the rest of the qubit pairs from the ten selected qubits are detailed in A. For idle-idle experiment, we apply exactly the same RB circuit simultaneously to the adjacent pair of qubits. While the duration of the idle period remains the same for both qubit pairs, the starting points of these idle periods differ slightly due to the differences in duration of the two neighboring CX pairs. We use the As Late As Possible policy to schedule the circuits. For the idle-driven experiment, while one qubit pair undergoes the RB circuit, its adjacent pair is subjected to a circuit composed of different random SU(4) blocks, which continuously drives crosstalk during the RB experiment, as shown in figure 3e. The number of SU(4)s in the circuit is determined to ensure it has the same duration as the RB circuit. This experimental arrangement ensures the simultaneous execution of both circuits. Since no idle time is inserted between SU(4)s, during the idle periods of the RB circuit, the neighboring qubits in the SU(4) circuit remain active thereby contributing to driven crosstalk. In the two experiments, the CX gate is aligned with the native direction of either ECR or DCX gates, determined in calibration to that of higher fidelity.
@@ III.2 Idle-idle Experiment - Simultaneous Randomized Benchmarking
The RB and simultaneous RB (SRB) results for the qubit pair ($q_{11},q_{14}$) are illustrated in figure 4, showcasing the comparison between various X2 sequences (X2pm, X2pm-stag, X2pm-stag-inv) against the free evolution. In the SRB experiment, it involves the adjacent qubit pair of ($q_{11},q_{14}$), which is ($q_{12}$, $q_{13}$), with $q_{14}$ and $q_{13}$ being the target qubits in their respective pairs and neighboring each other. The figure shows the connectivity of the two qubit pairs, detailing the type of CX gate used (ECR or DCX) and identifying whether a qubit acts as a control (C) or target (T) for the neighboring qubit connections between pairs. The qubit pair ($q_{11},q_{14}$) is highlighted in gray to indicate the specific results corresponding to this pair in the figure. The insertion of default X2 pulses results in unexpected behavior, even degrading the performance below that of the free evolution. The reason for this outcome is elaborated in B.
We also demonstrate the results for XY4 and XY8 sequences in figure 5. Given the negligible difference between the results for staggered DD and its inverse, only the staggered DD is depicted. We observe “dropouts” in fidelity for a short idle times in the XY8 sequence which can be explained by constructive interference of higher frequency spectral noise components [45, 46]. In addition, the XY8 sequence shows consistently worse results than XY4, which might be due to the accumulating single-qubit gate errors. Therefore, our subsequent discussions will focus primarily on the X2 and XY4 sequences.
Figure 4: Idle-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with different DD sequences: X2pm, staggered X2pm, and inverse of staggered X2pm. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when RB is applied to both qubit pairs simultaneously (SRB experiment). The significant difference in fidelity during free evolution between the RB and SRB experiments indicates additional crosstalk generated by the neighboring qubit pair. Note that “T” indicates the qubit serves as the target for the CX gate applied to the qubit pair. The results correspond to the gray qubit pair.
Figure 5: Idle-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with XY4pm, staggered XY4pm, XY8pm, and staggered XY8pm. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when RB is applied to both qubit pairs simultaneously (SRB experiment). The results correspond to the gray qubit pair.
@@ III.3 Driven-idle Experiment - Randomized Benchmarking and Random SU(4)
Figure 6: Driven-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when only RB is applied. (b) The same probability when of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair while SU(4) is applied to its adjacent qubit pair $(q_{12},q_{13})$. The results correspond to the gray qubit pair.
@@ IV Conclusion
Our proposed staggered DD protocol can be considered as an entry point for integrating DD sequences into multi-qubit quantum systems. Throughout the experiments, we have tailored our staggered DD protocol to fit X2, XY4, and XY8 sequences, incorporating them during unified idle times. Several advanced DD sequences, such as UDD [13], UR [14], concatenated DD (CDD) [48], and composite pulses [49], offer opportunities for further exploration. One natural following step is to apply the staggered DD protocol to these sequences and delve into DD scheduling methods for non-unified idle time in large quantum systems. Recently, many efforts are focused on understanding noise models for quantum hardware and developing corresponding error suppression or mitigation techniques [4]. An interesting direction is to synergize the staggered DD protocol with these techniques to enhance circuit fidelity further. Moreover, standard DD sequences have been employed to protect idle qubits for quantum error correction or detection codes [50, 51, 16, 52]. Incorporating our staggered DD protocol into quantum error correction codes offers a pathway toward achieving fault-tolerant quantum computing. Despite the positive impact of DD sequences, challenges such as dropouts or instability across different qubits persist. Additional experiments are necessary to deepen the understanding of hardware noise sources and to devise more hardware-oriented DD sequences.
Data Availability
The data that support the findings of this study will be openly available following an embargo at the following URL/DOI:
https://github.com/peachnuts/CrosstalkDD.
@@ Acknowledgements.
The authors acknowledge the IBM Quantum Network Hub at QuantUM Initiative of the Region Occitanie and University of Montpellier for providing additional quantum resources through its partnership with IBM. The authors would like to thank Bibek Pokharel for a careful reading of the manuscript and Ali Javadi-Abhari, Petar Jurcevic, and Christopher J. Wood for insightful discussions.
@@ Appendix A Various Experimental Configurations
There are two variants of CX gates in IBM quantum hardware: Echoed Cross-Resonance CX (ECR), and Direct CX (DCX). We select several qubit pairs from ibm_cairo including both ECR and DCX types of CNOTs. We list the properties of these qubits, including relaxation, decoherence, and error rate in table 2, and list the amplitudes, duration, and error rates of the qubit pairs in table 3. Note that, CX gate is natively uni-direction in IBM quantum hardware, and the opposite direction is realized by adding extra single qubit gates. Therefore, we show explicitly the control and target qubits for each qubit pair in the table. Generally, DCX gate exhibits a lower error rate and shorter duration compared with ECR gate. However, DCX gate is more susceptible to crosstalk, as illustrated in figure 7. Here, the CNOT error rate, derived from SRB, exhibits a notable increase compared to the error derived from independent RB for DCX, whereas the CNOT error rate for the ECR only experiences a slight increase.
Figure 7: We perform RB and SRB on two qubit pairs and calculate the corresponding CNOT error rate. The selected qubit pairs employ either ECR or DCX type of CX gate. An increase in the CNOT error rate observed during SRB as compared to RB indicates the impact of crosstalk.
@@
Table 2: Single-qubit properties including relaxation ($T_{1}$), decoherence ($T_{2}$), error rate determined by randomized benchmarking, frequency, and anharmonicity.
Qubit $T_{1}$ ($\mu s$) $T_{2}$ ($\mu s$) Error rate (%) Frequency (GHz) Anharmonicity (GHz)
0 73.69 144.16 0.02 5.26 0.32
1 149.91 137.63 0.02 5.4 0.34
4 71.49 29.85 0.04 5.19 0.34
7 104.08 68.32 0.08 5.05 0.34
10 102.86 26.56 0.02 5.23 0.34
11 78.89 9.8 0.04 5.13 0.34
12 126.13 214.87 0.01 5.11 0.34
13 115.74 148.44 0.02 5.28 0.34
14 105.07 209.43 0.02 5.04 0.34
15 124.46 164.34 0.01 4.96 0.34
Table 3: Two-qubit properties including type of entangling pulse for constructing CX, amplitude/duration of the pulses, error rate determined by randomized benchmarking, coupling rate $J$, and static $ZZ$ crosstalk strength.
CX(control-target) Type Amplitude Duration (dt) Error rate (%) $J$ (MHz) $ZZ$ (kHz)
(1, 0) ECR 0.91 2400 2.57 1.54 40.97
(4, 7) ECR 0.58 1184 1.64 2.05 59.04
(10, 12) DCX 0.42 1008 0.66 2.06 56.83
(11, 14) DCX 0.49 848 1.37 1.93 46.79
(12, 13) ECR 0.92 1760 1.4 2.06 66.12
(13, 14) DCX 0.33 992 0.49 2.11 103.04
(15, 12) ECR 0.52 2208 0.9 1.87 51.19
@@
are shown in (a) and (b), on $[(q_{10},q_{12})-(q_{13},q_{14})]$ are shown in (c) and (d), on $[(q_{12},q_{13})-(q_{11},q_{14})]$ are shown in (e) and (f), and on $[(q_{15},q_{12})-(q_{13},q_{14})]$ are shown in (g) and (h). The results correspond to the gray qubit pair.
Appendix B X2 Analysis
The exploration of circuit fidelity for both idle-idle and driven-idle experiments reveal that the insertion of X2 sequence leads to worse outcomes than the baseline, as shown in figure 9, which is potentially due to the imperfections in the X gate. However, this was not reproducible in simulations with an error model informed by quantum process tomography (QPT). The QPT result on the X gate on ibm_cairo shows a high process fidelity of 0.98. Since a static $ZZ$ interaction is always present between neighboring qubits, inserting an X gate introduces an $IY$ component due to the non-commutation between $ZZ$ and $IX$. The second X gate would typically amplify this $IY$ error. However, if we apply the second X gate in the opposite direction, denoted as Xm, the $IY$ component is introduced in the opposite direction and tends to cancel out the previous $IY$ error.